# An algorithm for forensic toolmark comparisons

### v2

@@ 6.1 Cross-validated classification performance
How well does classification work if we use the intersection point for the KM and KNM densities from experiment 1 as a threshold to classify between different-source and same-source pairs? Using the data from experiment 1, we use cross-validation to evaluate the classification performance. Specifically, we do the following: (1) Split the data into two folds, one with all the marks made by the tools marked by even numbers and the other with all the marks made by tools with odd numbers. (2) Train the algorithm using the marks from experiment 1 and obtain a threshold. (3) Use this threshold to classify between same-source and different-source pairs in the testing set. (4) Calculate the sensitivity and specificity of our algorithm by comparing the predicted outcome with the ground truth (KM or KNM). (5) Repeat this procedure with the second fold. (6) Average the sensitivity and specificity to obtain the cross-validated performance metrics.
@@
Table 3 shows the cross-validated sensitivity and specificity of our classification procedure. The table shows that the in-sample performance when using the threshold from experiment 1 (0.67) is very high. When using only the data from 3, the sensitivity is lower. It is difficult to say whether the lower sensitivity is due to having less data. However, after qualitatively observing the orange squares in the off-diagonals in Figures 7 and 8, it is not surprising to see that the classification is lower. Nevertheless, the performances are generally high for all three experiments.
@@ 6.2 Performance as a function of length
A specificity of 0.9, which corresponds to a 10% false positive rate, is reached at a signal length of about 1.5 mm. Once the signal length decreases further, there is not enough information to determine its source with any reasonable accuracy. At a signal length of just under 1 mm specificity drops to 50%, i.e. the accuracy of assessing whether a signal of that length matches a specific source is equal to a coin flip.
We see a similar dependence of accuracy on signal length in Hare et al. (2017b): at 37.5% of a land engraved area (corresponding to about .$375\times 2.2$ mm = 0.825 mm) specificity drops below 0.9 (for a sensitivity of about 0.7).
@@ 7 Discussion
Since our screwdrivers were consecutively manufactured, any sub-class characteristics generated by the manufacturing process, i.e., marks that were common among screwdrivers despite the fact that they are different tools, are likely included in our data. This means that our results hold not just for tools of a certain type of screwdriver, but even for tools made in the same batch, which could be difficult to differentiate.
@@
In future work, it would be useful to study how generating toolmarks at smaller angles of attack (i.e., less than 60) affects the classification performance, as well as other analyses that we presented with data from experiment 2. Baiker et al. (2014) had shown that smaller differences in angle did not affect classification, and our results agree with this, but that larger differences did affect classification. It would also be interesting to study the difference between making toolmaks with large and small screwdrivers. In our experimental setup, our mechanical rig only allowed us to make toolmarks at different angles with large screwdrivers, and toolmarks in different directions with small screwdrivers. Thus, it would be useful to have a complete factorial design, where toolmarks are generated wiht a different mechanical rig. Furthermore, it would be interesting to include degree of force applied as one of the factors, and how the factors interact with each other. We do not know how factors, such as angle of rotation, interaction of one tool with another as in wire cutters, tools that have strong class characteristics like serrated knives affect the classification performance of the algorithm out-of-sample.
@@
Although commonly used in forensics, classification by using a threshold from the density plots can be not a great idea because the threshold relies on parametric assumptions, it puts lots of weight on the tails of the distributions where there is the least amount of data available, and it is not clear how much the model performance depends on the training data. Would it help to diversify the training data for classification? And in what way? Would it be better to include toolmarks from a variety of different tools in the same training data, in a “kitchen sink” style? We did include the data from all three experiments in training densities, but we did not include them in this article because since experiments 2 and 3 have so much less data than experiment 1, the results did not change much from those for experiment 1. These questions remain for future research. It would be worthwhile to study the generalizability of this threshold to other tools and factors. Some methods have been trained on specific screwdrivers, there is evidence that they could be used to compare other tools as well (Baiker et al., 2014). This is promising because it means that, for toolmark comparisons to be accurate and useful, it might only be necessary to train algorithms using a small number of tools. Collecting data from all the tools that could be used to commit a crime is practically impossible, especially as the number and types of tools grows over time. Further research is necessary to determine how much each type of tool generalizes to other types of tools.
@@ 8 Conclusion
The shift from subjective to objective comparison methods in pattern-matching forensic disciplines has the potential to improve consistency, allow for the demonstration of process validation, and have less uncertainty in conclusions, all of which can reduce errors in comparisons and, therefore, improve the criminal justice system.

### v3

@@ 6.1 Cross-validated classification performance
How well does classification work if we use the intersection point for the KM and KNM densities from experiment 1 as a threshold to classify between different-source and same-source pairs? Using the data from experiment 1, we use cross-validation to evaluate the classification performance. Specifically, we do the following: (1) Split the data into two folds, one with all the marks made by the tools marked by even numbers and the other with all the marks made by tools with odd numbers. (2) Calculate the threshold using half the data (we call this the “training” set). (3) Use this threshold to classify between same-source and different-source pairs in the remaining data (the “testing” set. (4) Calculate the sensitivity and specificity of our algorithm by comparing the predicted outcome with the ground truth (KM or KNM). (5) Repeat this procedure with the second fold. (6) Average the sensitivity and specificity to obtain the cross-validated performance metrics.
@@
Table 3 shows the cross-validated sensitivity and specificity of our classification procedure. The table shows that the in-sample performance when using the threshold from experiment 1 (0.67) is very high. When using only the data from 3, the sensitivity is lower. It is difficult to say whether the lower sensitivity is due to having less data. However, after qualitatively observing the orange squares in the off-diagonals in Figures 8 and 9, it is not surprising to see that the classification is lower. Nevertheless, the performances are generally high for all three experiments.
Figure 14 shows the receiver operating characteristic (ROC) curve, which illustrates the performance of the binary classifier model at varying threshold values for the intersection of the KM and KNM curves. The experiment 1 threshold method is better at classifying, which makes sense since it is based on more data. Again, we do not recommend using experiments 2 and 3 to classify because they are based on too few data points.
Figure 14: Receiver operating characteristic (ROC) curve for the classifiation method based one the threshold of the KM and KNM density intersection for the three experiments.
@@ 6.2 Performance as a function of length
A specificity of 0.9, which corresponds to a 10% false positive rate, is reached at a signal length of about 1.5 mm. Once the signal length decreases further, there is not enough information to determine its source with any reasonable accuracy. At a signal length of just under 1 mm specificity drops to 50%, i.e. the accuracy of assessing whether a signal of that length matches a specific source is equal to a coin flip. Note that reducing the signal length could have caused FP errors in the location of the shorter segment due to alignment (i.e., registration). Studying the location of where the shorter segments found high CCF would be a useful exercise.
There is a similar dependence of accuracy on signal length in Hare et al. (2017b). The authors find that at 37.5% of a land engraved area (corresponding to about .$375\times 2.2$ mm = 0.825 mm) specificity drops below 0.9 (for a sensitivity of about 0.7).
@@ 7 Discussion
Our method holds for slotted screwdrivers and for screwdrivers that are made with a similar production method. To use this method to classify different types of tools, it may be necessary to perform similar experiments with different types of tools. A large database of 3D toolmarks could be created, and the more tools that researchers add, the better the method will be at classifying same- and different-source marks. Future research is needed to determine where the limits are of classifying using our training data.
@@
Since our screwdrivers were consecutively manufactured, any sub-class characteristics generated by the manufacturing process, i.e., marks that were common among screwdrivers despite the fact that they are different tools, are likely included in our data.
@@
In future work, it would be useful to study how generating toolmarks at smaller angles of attack (i.e., less than 60) affects the classification performance, as well as other analyses that we presented with data from experiment 2. Baiker et al. (2014) had shown that smaller differences in angle did not affect classification, and our results agree with this, but that larger differences did affect classification. It would also be interesting to study the difference between making toolmaks with large and small screwdrivers. In our experimental setup, our mechanical rig only allowed us to make toolmarks at different angles with large screwdrivers, and toolmarks in different directions with small screwdrivers. Thus, it would be useful to have a complete factorial design, where toolmarks are generated wiht a different mechanical rig. Furthermore, it would be interesting to include degree of force applied as one of the factors, and how the factors interact with each other. We do not know how factors, such as angle of rotation, interaction of one tool with another as in wire cutters, tools that have strong class characteristics like serrated knives affect the classification performance of the algorithm out-of-sample. This study was conducted using lead, as lead was the material we found to capture marks of the entire screwdriver tip. This is a a foundational study. It would be interesting to study how the marks change with different substrate materials, in addition to simply exhibiting marks of the incomplete screwdriver tip. It would also be interesting to study the effects on classification due to degradation of the tool with use and time.
@@
Although commonly used in forensics, classification by using a threshold from the density plots can be not a great idea because the threshold relies on parametric assumptions, it gives great importance to the tails of the distributions where there is the least amount of data available, and it is not clear how much the model performance depends on the training data. Would it help to diversify the training data for classification? And in what way? Would it be better to include toolmarks from a variety of different tools in the same training data, in a “kitchen sink” style? We did include the data from all three experiments in training densities, but we did not include them in this article because since experiments 2 and 3 have so much less data than experiment 1, the results did not change much from those for experiment 1. These questions remain for future research. It would be worthwhile to study the generalizability of this threshold to other tools and factors. Some methods have been trained on specific screwdrivers, there is evidence that they could be used to compare other tools as well (Baiker et al., 2014). This is promising because it means that, for toolmark comparisons to be accurate and useful, it might only be necessary to train algorithms using a small number of tools. Collecting data from all the tools that could be used to commit a crime is practically impossible, especially as the number and types of tools grows over time. Further research is necessary to determine how much each type of tool generalizes to other types of tools.
@@ 8 Conclusion
The shift from subjective to objective comparison methods in pattern-matching forensic disciplines has the potential to improve consistency, allow for the demonstration of process validation, and allow for more transparency and more possibilities for validation. All of these can reduce errors in comparisons and, therefore, improve the criminal justice system.